Write the converse of this statement.
If 3 - 2x = 13, then x = -5.

Answer by ikleyn(52894)

.
Write the converse of this statement.
If 3 - 2x = 13, then x = -5.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Wikipedia says:
      For the implication P → Q, the converse is Q → P

So, for the statement
     If 3 - 2x = 13, then x = -5

the converse statement is
     If x = -5, then 3 - 2x = 13